OpenStudy (anonymous):

A line contains the points (-2, 4) and (5, 4). Its equation is..... A. y = x + 7 B. y = x + 2 C. y = x + 5 D. y = 4 E. Cannot be determined

OpenStudy (texaschic101):

find the slope first...can you do that ?

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

sometimes

OpenStudy (texaschic101):

slope(m) = (y2 - y1) / (x2 - x1) (-2,4) -- x1 = -2, y1 = 4 (5,4) -- (x2 = 5, y2 = 4) lets sub the stuff in m = (4 - 4) / (5 - (-2) m = 0 / 7 The slope is 0, so it is a horizontal line. Therefore the equation is y = 4.

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

no, but thanks that helped

OpenStudy (texaschic101):

YW :)